3) x + 4 = 3 4) V8x - 1 = 3 X+1
Answer #1

la x + 4 3. X+1 L 2002 to X ( x + 1) + 4 roza X+) xta +4 - 3 xfl 22 +2 +4 = 3(441) + x2 +x+4= 32 +3 x²+x+4 - 32-3=0 = Esta on